Lectures by the Fields Medalists and by the Rolf Nevanlinna Prize Winner

The Fields Medalists Borcherds, Gowers and Kontsevich gave additional lectures in the afternoon program. They agreed that manuscripts of their talks should be included in the Proceedings. But finally, Maxim Kontsevich decided not to submit a manuscript of his lecture on “Motivic Galois Group and Deformation Quantization”, due to very recent developments in his field of interest.

Richard E. Borcherds: What is Moonshine? . . . . . . . . . . . . . . . . . . . . . I 607
W. T. Gowers: Fourier Analysis and Szemerédi’s Theorem . . . . . . . . . I 617

Curtis T. McMullen: Rigidity and Inflexibility in Conformal Dynamics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 841

see also: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . I 630
Peter W. Shor: Quantum Computing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . I 467

Documenta Mathematica · Extra Volume ICM 1998 · I

606

Documenta Mathematica · Extra Volume ICM 1998 · I